OpenStudy (bbypanda16):

Help please? if JKhas endpoints J(-2,3) and k(6,5), and LN has endpoints L(0,7) and N(4,1), then JK and LN have the same midpoint. If GH has endpoints G(-8,1) and H(4,5), then the midpoint M of GH lies on the line y= -x+1.

OpenStudy (bbypanda16):

Will fan and medal!!!

OpenStudy (danjs):

ok, what is the question

OpenStudy (bbypanda16):

if JKhas endpoints J(-2,3) and k(6,5), and LN has endpoints L(0,7) and N(4,1), then JK and LN have the same midpoint. @DanJS

OpenStudy (danjs):

ok do you have to just show that? It is not asking anything

OpenStudy (bbypanda16):

the question says show thateach question is rue

OpenStudy (bbypanda16):

true*

OpenStudy (danjs):

The midpoint of a line segment is the average of the x and average of the y values.. two endpoints (x1, y1) and (x2, y2) \[\large (\frac{ x1+x2 }{ 2 }~~~~,~~~~\frac{ y1+y2 }{ 2 })\]

OpenStudy (danjs):

Do that for both the segments to show they have the same midpoint value

OpenStudy (bbypanda16):

okay, than you ::)

OpenStudy (danjs):

maybe make a graph too, welcome

OpenStudy (danjs):

If GH has endpoints G(-8,1) and H(4,5), then the midpoint M of GH lies on the line y= -x+1. This you use the midpoint in the line equation and show that it is true... midpoint - (x , y) plug x and y into line, should be true